News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 10 June 2011 9.00am EDT

During the last week the observing schedule was interrupted by a high radiation event that stopped the daily load at 11:23am EDT on Jun 7 and activated the science instrument safing sequence (SCS 107). All spacecraft actions were nominal. A real-time procedure was executed on Jun 7 to re-enable the spacecraft momentum monitor, which had been left disabled by the stoppage of the daily load. A real-time procedure was executed on Jun 8 to verify that the radiation environment as observed by the HRC anticoincidence shield had decreased to an acceptable level and a replanned schedule was started. Several Norma Region observations were impacted and will be rescheduled. The replanned schedule included an observation of PTF11eon, which was accepted as a Target of Opportunity on Jun 3.

Chandra passed through the 21st, 22nd, and 23rd eclipses of the spring season on Jun 4, 7, and 9 respectively. The 21st and 22nd eclipses proceeded with nominal power and thermal performance. Real-time procedures were executed to disable SCS-29 and to set the short-eclipse flag to false (medium) on Jun 8 following the completion of the planned eclipse season activities. The 23rd and final eclipse of the season was a penumbra only type and was expected to require no action. However, the thermal trend on the arrays during this event resulted in an unexpected eclipse detection and connection of the batteries to the spacecraft bus. The handling of this event by the on-board EPS manager software went nominally.

A real-time procedure was executed on Jun 9 to update the on-board ephemeris.

The schedule of targets for the next week is shown below

------------------------------------------
Radiation Belts                     Jun 12
HD50806                ACIS-S
NormaRegion (8 obs)    ACIS-I       Jun 13
Radiation Belts                     Jun 15
PSRJ0726-2612          ACIS-S
NormaRegion (8 obs)    ACIS-I
Radiation Belts                     Jun 17
1ES1553+113            HRC-S/LETG   Jun 18

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports